OpenStudy (anonymous):

what is the solution set |x-8| > 5.6 and how do i go about it for future problems?

OpenStudy (anonymous):

its asking for interval notation.

OpenStudy (dumbcow):

i got my inequalities backwards

OpenStudy (anonymous):

so what exactly does that mean? lol

OpenStudy (dumbcow):

5.6 <x+8 <-5.6

OpenStudy (anonymous):

so i am getting -2.4<x<-13.6 ?

OpenStudy (anonymous):

|x-8| means the modulus value (or positive value) of x-8 For example if x=4, x-8 = (4)-8 = -4 However, |x-8| = 4 as you take the positive value. \[Therefore |x-8|>5.6 is the same as x-8>5.6 and x-8<-5.6\] So your answers are x>13.6 and x<3.6. Understand?

OpenStudy (anonymous):

x<2.4 sorry

OpenStudy (anonymous):

so there is no 13.6 involved?

OpenStudy (anonymous):

\[x \in(-\infty,2.4)\cup(13.6,\infty)\]

OpenStudy (anonymous):

x>13.6 and x<2.4

OpenStudy (anonymous):

ahh ok! thank you!